# MetaOOD: Automatic Selection of OOD Detection Models

### v1

@@ Abstract
How can we automatically select an out-of-distribution (OOD) detection model for various underlying tasks? This is crucial for maintaining the reliability of open-world applications by identifying data distribution shifts, particularly in critical domains such as online transactions, autonomous driving, and real-time patient diagnosis. Despite the availability of numerous OOD detection methods, the challenge of selecting an optimal model for diverse tasks remains largely underexplored, especially in scenarios lacking ground truth labels. In this work, we introduce MetaOOD, the first zero-shot, unsupervised framework that utilizes meta-learning to select an OOD detection model automatically. As a meta-learning approach, MetaOOD leverages historical performance data of existing methods across various benchmark OOD datasets, enabling the effective selection of a suitable model for new datasets without the need for labeled data at the test time. To quantify task similarities more accurately, we introduce language model-based embeddings that capture the distinctive OOD characteristics of both datasets and detection models. Through extensive experimentation with 24 unique test dataset pairs to choose from among 11 OOD detection models, we demonstrate that the MetaOOD significantly outperforms existing methods and only brings marginal time overhead. Our results, validated by Wilcoxon statistical tests, show that MetaOOD surpasses a diverse group of 11 baselines, including established OOD detectors and advanced unsupervised selection methods.
@@ 1 Introduction
Out-of-distribution (OOD) detection is the process of identifying data points that deviate significantly from the distribution of the training data. This capability is essential for ensuring the reliability of machine learning (ML) models when they encounter new, unseen data (Yang et al., 2021). Common applications of OOD detection include safety-critical systems like autonomous vehicles (Filos et al., 2020) and medical diagnosis (Ulmer et al., 2020) to prevent erroneous predictions. Notably, many OOD detection algorithms are unsupervised due to the high annotation cost, leveraging statistical methods or reconstruction errors from models such as autoencoders to identify deviated samples without labeled OOD data (Yang et al., 2022; Dong et al., 2024). However, despite the variety of OOD detection algorithms, each adepts at identifying different aspects of in-distribution (ID) and OOD data, there lacks a systematic method for choosing the best OOD detection model(s) under the unsupervised setting without labels. It is acknowledged that each OOD detection algorithm might excel in specific scenarios but may not perform well universally due to the no-free-lunch theorem (Wolpert & Macready, 1997). Moreover, without labels, it is difficult to objectively evaluate and compare the performance of different OOD detection models.
@@
Present Work. In the most related field, unsupervised outlier detection (OD) has benefitted significantly from meta-learning, with models like MetaOD (Zhao et al., 2021), ELECT (Zhao et al., 2022a), ADGym (Jiang et al., 2024), and HyPer (Ding et al., 2024) demonstrating notable advancements. By leveraging historical performance data, these methods estimate a model’s efficacy on new datasets. However, they are not directly adaptable to OOD detection due to several key differences: First, OD and OOD detection differ in their problem settings. OD models are trained on both normal and outlier samples (Zhao et al., 2019; Ma et al., 2023), focusing on anomalies within a single distribution (Zhao et al., 2022b; Zhao, 2024). In contrast, OOD detection involves training solely on in-distribution (ID) data and aims to identify samples from entirely different distributions, often across multiple datasets. Second, OOD detection tasks primarily deal images, presenting greater complexity compared to time series and tabular data in OD. Third, the embeddings used to measure dataset similarity in OD do not translate well to OOD contexts. Current OD embedding generation are largely heuristic and feature-specific (Zhao et al., 2021), inadequate for the diverse and complex nature of OOD datasets, especially when considering the different data modalities. Thus, addressing these challenges with a tailored method for OOD model selection is crucial.
@@
Figure 1: MetaOOD overview (§3.2); offline meta-training phase is shown on the top (§3.3)–the key is to train a meta performance predictor $f$ (denoted in ) to map language embeddings of the datasets and models to their performance $\mathbf{P}$; the online model selection (§3.4) is shown at the bottom by transferring the meta-predictor $f$ to predict the test data paired with OOD models for selection.
@@
Our Work. In this study, we introduce MetaOOD, the first unsupervised OOD  [^1] For brevity, we may refer OOD detection as OOD throughout the paper. model selection method that employs meta-learning. Figure 1 illustrates the overall workflow of the proposed approach. The key rationale is that an OOD model is likely to perform well on a new dataset if it excels in a similar historical dataset. As a meta-learning approach, we train a suite of OOD detection models offline using a variety of carefully curated datasets to gauge their performance across different scenarios during the meta-train phase (Figure 1, top). When a new dataset arrives, we apply knowledge derived from historical data to select an appropriate OOD model (Figure 1, bottom). This selection is based on the similarity between the new dataset and those used in the meta-train phase. To enhance the accuracy of this similarity assessment, we designed two versions of data embeddings via specialized OOD meta-features and language embeddings from language models. As our experiment shows, these language model-based embeddings effectively capture complex, nuanced dataset characteristics that traditional meta-features might miss. We summarize our technical contributions:
@@
  • Specialized Embeddings for OOD Datasets and Models. We use language model-generated features to quantify the similarity among OOD tasks, facilitating a better understanding of OOD characteristics and enhancing OOD detection model selection.
@@
  • Accessibility and Reproducibility. We release the testbed, corresponding code, and the proposed meta-learner at https://anonymous.4open.science/r/metaood-DA67.
@@ 2 Related Work
2.1 Unsupervised OOD Model Selection
OOD detection faces the challenge of both OOD samples and OOD distributions being unknown during the training phase (Hendrycks et al., 2019; Liang et al., 2018). Consequently, selecting models for OOD relies solely on in-distribution or training data (Lee et al., 2018b). Once deployed in an open-world setting, the OOD detector often handles a diverse range of test inputs originating from various OOD distributions (Yang et al., 2021), and consequently, unsupervised methods are expected to be more suitable for selecting OOD detection models (Liu et al., 2020b). Within these unsupervised OOD detection models, some recent works focus on actively choosing sample-wise detectors (e.g., choosing different groups of detectors for different input samples) (Xue et al., 2024), while it is not doing actual model selection at the single model level as this work tries to address.
Most existing methods rely on trial-and-error or empirical heuristics. For instance, the simplest approach may be not to actively “select” a model, but just use popular OOD detection models like Maximum Softmax (Hendrycks & Gimpel, 2017) and ODIN (Liang et al., 2017). There are other simple approaches; the confidence level of the ID data may be used as an indicator for OOD detection model selection. However, such a simple method may lead to a low true positive rate (TPR) (Hendrycks & Gimpel, 2017). Another direction is similarity-based methods, where the model selection is based on the similarity or cluster of the dataset, which has also been utilized for algorithm recommendation (Kadioglu et al., 2010; Nikolić et al., 2013; Xu et al., 2012; Misir & Sebag, 2017). We thus include these algorithms as our baselines in §4.
2.2 Supervised OOD Model Selection
There has been considerable research into the use of supervised model selection techniques for effectively training and evaluating predictive models on labeled datasets (James et al., 2013; Hastie et al., 2015; Tibshirani, 1996). These methods are particularly valuable in scenarios where generalization to unseen data is critical. Model selection in supervised learning encompasses a variety of strategies, including cross-validation, grid search, and performance metrics such as accuracy and F1 score to assess model efficacy. Randomized (Bergstra & Bengio, 2012), bandit-based (Li et al., 2017), and Bayesian optimization (BO) techniques (Shahriari et al., 2015) are various SOTA approaches to hyperparameter optimization and/or algorithm selection. However, it is notable that such methods do not apply to unsupervised OOD model selection, as ground truth values are absent.
@@ 3.2 Problem Statement and Framework Overview
Given a new, unseen dataset pair for OOD detection, the goal is to select the best model from a heterogeneous set of OOD detection models without requiring any model evaluations at the test.
In this work, we leverage meta-learning to transfer model performance information from prior experiences to the new OOD task. Meta-learning (Vanschoren, 2018), often called “learning to learn,” is a technique where an algorithm learns from a collection of historical/meta tasks and uses this experience to perform well on new, unseen tasks. The rationale is that an OOD model is likely to outperform on a new dataset if it excels in a similar historical dataset. This approach is particularly useful when evaluation is infeasible or costly due to a lack of labels or the need for rapid adaptation.
@@
  • A collection of $n$ historical (i.e., meta-train) OOD detection dataset pairs, $\mathcal{D}_{\text{train}}=\{D_{1},\ldots,D_{n}\}$ with ground truth labels, i.e., ${D}=\{\mathbf{X}_{\text{train}},(\mathbf{X}_{\text{test}},\underline{\mathbf{y}_{\text{test}}})\}$
  • Historical performance $\mathbf{P}$ of the pre-set model set =$\mathcal{M}=\{M_{1},\ldots,M_{m}\}$ (with $m$ models), on the meta-train datasets. We refer to $\mathbf{P}\in\mathbb{R}^{n\times m}$ as the performance matrix, where $\mathbf{P}_{i,j}$ corresponds to the $j$-th model $M_{j}$’s performance on the $i$-th meta-train dataset $D_{i}$.
@@
In a nutshell, our MetaOOD consists of two phases: (i) offline (meta-) training of the meta-learner on $\mathcal{D}_{\text{train}}$ where the goal is to learn the mapping from OOD detection models’ performance on various (historical/meta) datasets, and (ii) online model selection that uses the meta-train information to choose the model at test time for the new dataset $D_{\text{new}}$. Figure 1 outlines the workflow and key elements of MetaOOD, with the offline training phase shown in white and the online model selection stage shown in grey. The details of the two phases are discussed in §3.3 and §3.4, respectively.
@@ 3.3 Offline Meta-Training
During the offline training, we generate embeddings for both dataset pair $\mathcal{D}_{\text{train}}$ and method $\mathcal{M}$, and train the latent mapping from these embeddings to the performances $\mathbf{P}$. The meta-learner can generalize and select the best-performing model for new, unseen datasets by learning the relationship between $\{\mathcal{D}_{\text{train}},\mathcal{M}\}\rightarrow\mathbf{P}$. Prior works have shown that performance mapping is empirically learnable, although imperfect, in related fields like outlier detection (Zhao, 2024).
To predict the performance of the candidate model on a new dataset pair, we propose training a meta-predictor as a regression problem. The input to the meta-predictor consists of ${E_{i}^{\text{meta}},E_{j}^{\text{model}}}$, corresponding to the embedding of the $i$-th dataset pair and the embedding of the $j$-th OOD detector. Dataset embedding of dataset pair $\mathcal{D}$ is denoted as $E^{\text{data}}=\psi(\mathcal{D})$, and method embedding extracted from $\mathcal{M}$ is denoted as $E^{\text{model}}=\phi(\mathcal{M},M)$; we provide more details below on the embedding generation in §3.3.1. Our goal is to train the meta-predictor $f$ [^2] The format of $f$ can be any regression models; in this work, we use an XGBoost Chen & Guestrin (2016) model due to its balance of simplicity and expressiveness, as well as strong feature selection characteristic. to map the characteristics of the datasets and the OOD detectors to their corresponding performance ranking across all historical dataset pairs. The steps of the (offline) meta-train are shown in Fig. 1, top and Appx. Algo. 1.
@@ 3.3.1 Data and Model Embeddings
Data and model embeddings, as inputs of $f$, offer compact, standardized context about the data and models utilized in the meta-learning process. Instead of directly using data with arbitrary sizes, we hope to generate embeddings that can effectively represent the data and help develop models that can adapt rapidly to new tasks. In this work, we try two types of data and model embeddings: (1) classical (data) meta-features and (method) one-hot encoding embedding, and (2) language model based approach to get data and model embedding via language models.
@@
Statistical meta-features capture the underlying data distributions’ characteristics, including variance, skewness, covariance, and other relevant statistics of individual features and their combinations. For image datasets utilized in OOD, we incorporate meta-features that reflect pertinent image characteristics (Aguiar et al., 2019), such as: (i) Color-based: Simple statistical measures from color channels; (ii) Border-based: Statistical measures obtained after applying border-detector filters; (iii) Histogram: Statistics from histograms of color and intensity; and (iv) Texture features: Values derived from an image’s texture, analyzed using the co-occurrence matrix and Fast Fourier Transform.
Landmarker features summarize the performance of specific learning algorithms on a dataset, providing quick and effective estimates of algorithm performance. These features, derived from evaluating simple/fast models, offer a snapshot of dataset characteristics and approximate the performance of more sophisticated models in specific tasks. For OOD detection, our landmarkers focus on features related to Softmax probability outputs that can be rapidly generated without model fitting. Detailed descriptions and a complete list of our OOD meta-features are in Appx. Table  A.
@@
For OOD model embeddings, we take the textual descriptions of methods from the Pytorch-OOD library (Kirchheim et al., 2022) as input. This includes detailed information about the OOD detector, such as its components, whether it needs fitting, and what is used as OOD indicator. An example of this is the description for the Openmax method (Bendale & Boult, 2016):
@@
The complete lists of the datasets and methods used are listed in §4.1. In brief, this approach leverages language models’ natural language processing capabilities to transform qualitative descriptions into quantitative embeddings. Compared to the traditional meta-feature approach, it is more computationally efficient and has better generalization ability. Also, it is interesting to see whether we could use language models to choose an OOD model directly. We provide all these analyses in §4.

### v2

@@ Abstract
How can we automatically select an out-of-distribution (OOD) detection model for various underlying tasks? This is crucial for maintaining the reliability of open-world applications by identifying data distribution shifts, particularly in critical domains such as online transactions, autonomous driving, and real-time patient diagnosis. Despite the availability of numerous OOD detection methods, the challenge of selecting an optimal model for diverse tasks remains largely underexplored, especially in scenarios lacking ground truth labels. In this work, we introduce MetaOOD, the first zero-shot, unsupervised framework that utilizes meta-learning to select an OOD detection model automatically. As a meta-learning approach, MetaOOD leverages historical performance data of existing methods across various benchmark OOD detection datasets, enabling the effective selection of a suitable model for new datasets without the need for labeled data at the test time. To quantify task similarities more accurately, we introduce language model-based embeddings that capture the distinctive OOD characteristics of both datasets and detection models. Through extensive experimentation with 24 unique test dataset pairs to choose from among 11 OOD detection models, we demonstrate that MetaOOD significantly outperforms existing methods and only brings marginal time overhead. Our results, validated by Wilcoxon statistical tests, show that MetaOOD surpasses a diverse group of 11 baselines, including established OOD detectors and advanced unsupervised selection methods.
@@ 1 Introduction
Out-of-distribution (OOD) detection is the process of identifying data points that deviate significantly from the distribution of the training data. This capability is essential for ensuring the reliability of machine learning models when they encounter new, unseen data (Yang et al., 2021). Common applications of OOD detection include safety-critical systems like autonomous vehicles (Filos et al., 2020; Li et al., 2024a) and medical diagnosis (Ulmer et al., 2020) to prevent erroneous predictions. Notably, many OOD detection algorithms are unsupervised due to the high annotation cost, leveraging statistical methods or reconstruction errors from models such as autoencoders to identify deviated samples without labeled OOD data (Yang et al., 2022; Dong et al., 2024; Sun et al., 2022). However, despite the variety of OOD detection algorithms, each adepts at identifying different aspects of in-distribution (ID) and OOD data, there lacks a systematic method for choosing the best OOD detection model(s) under the unsupervised setting without labels. It is acknowledged that each OOD detection algorithm might excel in specific scenarios but may not perform well universally due to the no-free-lunch theorem (Wolpert & Macready, 1997). Moreover, without labels, it is difficult to objectively evaluate and compare the performance of different OOD detection models.
@@
Present Work. In the most related field, unsupervised outlier detection (OD) has benefitted significantly from meta-learning, with models like MetaOD (Zhao et al., 2021), ELECT (Zhao et al., 2022a), ADGym (Jiang et al., 2024), and HyPer (Ding et al., 2024) demonstrating notable advancements. By leveraging historical performance data, these methods estimate a model’s efficacy on new datasets. However, they are not directly adaptable to OOD detection due to several key differences: First, OD and OOD detection differ in their problem settings. OD models are trained on both normal and outlier samples (Zhao et al., 2019; Ma et al., 2023), focusing on anomalies within a single distribution (Zhao et al., 2022b; Zhao, 2024; Chen et al., 2024). In contrast, OOD detection involves training solely on ID data and aims to identify samples from entirely different distributions, often across multiple datasets (Li et al., 2024b; Liu et al., 2024). Second, OOD detection tasks primarily deal with images, presenting greater complexity compared to time series and tabular data in OD. Third, the embeddings used to measure dataset similarity in OD do not translate well to OOD detection contexts. Current OD embedding generation are largely heuristic and feature-specific (Zhao et al., 2021), inadequate for the diverse and complex nature of OOD datasets, especially when considering the different data modalities. Thus, addressing these challenges with a tailored method for OOD detection model selection is crucial.
@@
Figure 1: MetaOOD overview (§3.2); offline meta-training phase is shown on the top (§3.3)—the key is to train a meta performance predictor $f$ (denoted in ) to map language embeddings of the datasets and models to their performance $\mathbf{P}$; the online model selection (§3.4) is shown at the bottom by transferring the meta-predictor $f$ to predict the test data paired with OOD detectors for selection.
@@
Our Work. In this study, we introduce MetaOOD, the first unsupervised OOD detection model selection method that employs meta-learning. Fig. 1 illustrates the overall workflow of the proposed approach. The idea is that an OOD detector that performed well on similar historical datasets is likely to excel on new ones. As a meta-learning approach, we train a suite of OOD detection models offline using a variety of carefully curated datasets to gauge their performance across different scenarios during the meta-train phase (Fig.1, top). When a new dataset arrives, we apply knowledge derived from historical data to select an appropriate OOD detection model (Fig. 1, bottom). This selection is based on the similarity between the new dataset and those used in the meta-train phase. To enhance the accuracy of this similarity assessment, we designed two versions of data embeddings via specialized OOD meta-features and language embeddings from language models. As our experiment shows, these language model-based embeddings effectively capture complex, nuanced dataset characteristics that traditional meta-features might miss. We summarize our technical contributions:
@@
  • Specialized Embeddings for OOD Datasets and Models. We use language model-generated features to quantify the similarity among OOD detection tasks, facilitating a better understanding of OOD characteristics and enhancing OOD detection model selection.
@@
  • Accessibility and Reproducibility. We release the testbed, corresponding code, and the proposed meta-learner at https://github.com/yqin43/metaood.
@@ 2 Related Work
2.1 Unsupervised OOD detection Model Selection
OOD detection faces the challenge of both OOD samples and OOD distributions being unknown during the training phase (Hendrycks et al., 2019; Liang et al., 2018). Consequently, selecting models for OOD detection relies solely on in-distribution or training data (Lee et al., 2018b). Once deployed in an open-world setting, the OOD detector often handles a diverse range of test inputs originating from various OOD distributions (Yang et al., 2021), and consequently, unsupervised methods are expected to be more suitable for selecting OOD detection models (Liu et al., 2020b). Within these unsupervised methods, some recent works focus on actively choosing sample-wise detectors (e.g., choosing different groups of detectors for different input samples) (Xue et al., 2024), while it is not doing actual model selection at the single model level as this work tries to address.
Most existing methods rely on trial-and-error or empirical heuristics. For instance, the simplest approach may be not to actively “select” a model, but just use popular OOD detectors like Maximum Softmax (Hendrycks & Gimpel, 2017) and ODIN (Liang et al., 2017). There are other simple approaches; the confidence level of the ID data may be used as an indicator for OOD detection model selection. However, such a simple method may lead to a low true positive rate (TPR) (Hendrycks & Gimpel, 2017). Another direction is similarity-based methods, where the model selection is based on the similarity or cluster of the dataset, which has also been utilized for algorithm recommendation (Kadioglu et al., 2010; Nikolić et al., 2013; Xu et al., 2012; Misir & Sebag, 2017). We thus include these algorithms as our baselines in §4.
2.2 Supervised OOD detection Model Selection
There has been considerable research into the use of supervised model selection techniques for effectively training and evaluating predictive models on labeled datasets (James et al., 2013; Hastie et al., 2015; Tibshirani, 1996). These methods are particularly valuable in scenarios where generalization to unseen data is critical. Model selection in supervised learning encompasses a variety of strategies, including cross-validation, grid search, and performance metrics such as accuracy and F1 score to assess model efficacy. Randomized (Bergstra & Bengio, 2012), bandit-based (Li et al., 2017), and Bayesian optimization (BO) techniques (Shahriari et al., 2015) are various SOTA approaches to hyperparameter optimization and/or algorithm selection. However, it is notable that such methods do not apply to unsupervised OOD detection model selection, as ground truth values are absent.
@@ 3.2 Problem Statement and Framework Overview
Given a new, unseen dataset pair for OOD detection, the goal is to select the best model [^1] The use of the term ’models’ is more for following the tradition of model selection research. from a heterogeneous set of OOD detection models without requiring any model evaluations at test. In this work, we leverage meta-learning to transfer model performance information from prior experiences to the new OOD detection task. Meta-learning (Vanschoren, 2018), often called “learning to learn”, is a technique where an algorithm learns from a collection of historical/meta tasks and uses this experience to perform well on new, unseen tasks. The rationale is that an OOD detector is likely to outperform on a new dataset if it excels in a similar historical dataset. This approach is particularly useful when evaluation is infeasible or costly due to a lack of labels or the need for rapid adaptation.
@@
  • A collection of $n$ historical (i.e., meta-train) OOD detection dataset pairs, $\mathcal{D}_{\text{train}}=\{D_{1},\ldots,D_{n}\}$ with ground truth labels, i.e., ${D}=\{\mathbf{X}_{\text{train}},(\mathbf{X}_{\text{test}},\underline{\mathbf{y}_{\text{test}}})\}$.
  • Historical performance $\mathbf{P}$ of the pre-set model set =$\mathcal{M}=\{M_{1},\ldots,M_{m}\}$ (with $m$ models), on the meta-train datasets. We refer to $\mathbf{P}\in\mathbb{R}^{n\times m}$ as the performance matrix, where $\mathbf{P}_{i,j}$ corresponds to the $j$-th model $M_{j}$’s performance on the $i$-th meta-train dataset pair $D_{i}$.
@@
In a nutshell, our MetaOOD consists of two phases: (i) offline (meta-) training of the meta-learner on $\mathcal{D}_{\text{train}}$ where the goal is to learn the mapping from OOD detection models’ performance on various (historical/meta) datasets, and (ii) online model selection that uses the meta-train information to choose the model at test time for the new dataset $D_{\text{new}}$. Fig.1 outlines the workflow and key elements of MetaOOD, with the offline training phase shown in white and the online model selection stage shown in grey. The details of the two phases are discussed in §3.3 and §3.4, respectively.
@@ 3.3 Offline Meta-Training
During offline training, we generate embeddings for dataset pair $\mathcal{D}_{\text{train}}$ and method $\mathcal{M}$, and train the latent mapping from these embeddings to the performances $\mathbf{P}$. The meta-learner can generalize and select the best-performing model for new, unseen datasets by learning the relationship between $\{\mathcal{D}_{\text{train}},\mathcal{M}\}\rightarrow\mathbf{P}$. Note this training process is supervised. Prior works have shown that performance mapping is empirically learnable, although imperfect, in related fields like OD (Zhao, 2024).
To predict the performance of the candidate model on a new dataset pair, we propose training a meta-predictor as a regression problem. The input to the meta-predictor consists of ${E_{i}^{\text{meta}},E_{j}^{\text{model}}}$, corresponding to the embedding of the $i$-th dataset pair and the embedding of the $j$-th OOD detector. Dataset embedding of dataset pair $\mathcal{D}$ is denoted as $E^{\text{data}}=\psi(\mathcal{D})$, and method embedding extracted from $\mathcal{M}$ is denoted as $E^{\text{model}}=\phi(\mathcal{M},M)$; we provide more details below on the embedding generation in §3.3.1. Our goal is to train the meta-predictor $f$ [^2] The format of $f$ can be any regression models; in this work, we use an XGBoost (Chen & Guestrin, 2016) model due to its balance of simplicity and expressiveness, as well as strong feature selection characteristic. to map the characteristics of the datasets and the OOD detectors to their corresponding performance ranking across all historical dataset pairs. The steps of the (offline) meta-train are shown in Fig. 1, top and Appx. Algo. 1.
@@ 3.3.1 Data and Model Embeddings
Data and model embeddings, as inputs of $f$, offer compact, standardized context about the data and models utilized in the meta-learning process. Instead of directly using data with arbitrary sizes, we hope to generate embeddings that can effectively represent the data and help develop models that can adapt rapidly to new tasks. In this work, we try two types of data and model embeddings: (1) classical (data) meta-features and (method) one-hot encoding embedding, and (2) language model-based approach to get data and model embeddings via language models.
@@
Statistical meta-features capture the underlying data distributions’ characteristics, including variance, skewness, and other relevant statistics of individual features and their combinations. For image datasets used in OOD detection, we incorporate meta-features that reflect pertinent image characteristics (Aguiar et al., 2019), such as: (i) Color-based: Simple statistical measures from color channels; (ii) Border-based: Statistical measures obtained after applying border-detector filters; (iii) Histogram: Statistics from histograms of color and intensity; and (iv) Texture features: Values derived from an image’s texture, analyzed using the co-occurrence matrix and Fast Fourier Transform.
Landmarker features summarize the performance of specific learning algorithms on a dataset, providing quick and effective estimates of algorithm performance. These features, derived from evaluating simple/fast models, offer a snapshot of dataset characteristics and approximate the performance of more sophisticated models in specific tasks. For OOD detection, our landmarkers focus on features related to Softmax probability outputs that can be rapidly generated without model fitting. Detailed descriptions and a complete list of our OOD meta-features are in Appx. Tab. B.
@@
For OOD detection model embeddings, we take the textual descriptions of methods from the Pytorch-OOD library (Kirchheim et al., 2022) as input. This includes detailed information about the OOD detector, such as its components, whether it needs fitting, and what is used as OOD indicator. An example of this is the description for the Openmax method (Bendale & Boult, 2016):
@@
The complete lists of the datasets and methods used are listed in §4.1. In brief, this approach leverages language models’ natural language processing capabilities to transform qualitative descriptions into quantitative embeddings. Compared to the traditional meta-feature approach, it is more computationally efficient and has better generalization ability. Also, it is interesting to see whether we could use language models to choose an OOD detector directly. We provide all these analyses in §4.